(1/c+1/d) / (c/d-d/c)
A)1/c-d
B)1/c+d
C)c+d
D)c-d

Answer by al_prac(2)

Let us simplify the numerator of the fraction given:
1/c + 1/d = (d+c)/cd
Let us simplify the denominator of the fraction:
(c/d) - (d/c) = (c^2 - d^2) /cd
= ((c-d)(c+d))/cd
So now our fraction will be:
(((c+d)/cd))/(((c+d)(c-d))/cd)))) = 1/(c-d)
So answer is A, 1/(c-d).